What is the length of a simple pendulum used in a grandfather clock that has one second between it's tick and it's tock on Earth?

1
Expert's answer
2021-12-10T13:21:03-0500

"T=2\pi\sqrt{l/g}\to l=\frac{gT^2}{4\pi^2}=\frac{9.81\cdot1^2}{4\cdot3.14^2}=0.249\ (m)"